OpenStudy (enricoanders):

Find an equation for the line perpendicular to y=-17x+2 and goes through the point (-8,-10). Write your answer in the form y=mx+b.

OpenStudy (thatonegirl_):

To write an equation for a line, you need a slope and a point (x,y). For the slope: Perpendicular means that it will have an opposite reciprocal slope. (flip the numerator and denominator and make it positive instead of negative). So your new perpendicular slope will be 1/17. You already have a point (-8,-10). So now, plug in what you know in y=mx+b (y=-10, m=1/17 x=-8 and b we don't know so that's what you have to solve for). Once you find your b value, you can write y=1/17x+(b).

OpenStudy (enricoanders):

y=1/17x-2

OpenStudy (enricoanders):

@thatonegirl_